# DREAMS: Preserving both Local and Global Structure in Dimensionality Reduction

### v1

@@ 4.2 Linear decoding regularization
$\widetilde{W}=\underset{W\in\mathbb{R}^{m\times d}}{\arg\min}\|X-XWW^{T}\|_{2}^{2}\text{ subject to }W^{T}W=I_{d}.$ (4)
Since the PCA embedding is given by $\widetilde{Y}=X\widetilde{W}$, the DREAMS loss including the decoding regularizer becomes
$\mathcal{L}(Y,D)=(1-\lambda)\mathcal{L}_{t-\text{SNE}}+\lambda\left\|X-(YD^{T}+b)\right\|_{2}^{2},$
with $D\in\mathbb{R}^{m\times d}$ being a trainable linear decoder and $b\in\mathbb{R}^{m}$ a trainable bias term that is added row-wise and allows to handle uncentered embeddings. In this setup, by minimizing the reconstruction error, the decoder is responsible for the global structure preservation by pushing the embedding towards the PCA structure that has the minimal reconstruction error. If $D$ were constrained to be orthogonal $D^{T}D=I_{d}$, the optimum of the regularization term would be $D=\widetilde{W}$ (Plaut, 2018; Nazari et al., 2023). Although we do not explicitly enforce orthogonality, we observed that the learned linear decoder $D$ naturally tends to be approximately orthogonal.
@@
For comparison, we also implemented a version of DREAMS using the CNE backend with precomputed PCA regularization as in Section 4.1. We will refer to this version as DREAMS-CNE. Here the gradient with respect to $Y$ was computed using autodifferentiation, including the $\|Y\|$ contribution to $\alpha$.
@@ 5 Experimental setup
5.1 Datasets and metrics
Table 1: Datasets used in our experiments with a description of their local and global structure. For MNIST, we only used the training set.
@@ 5.1 Datasets and metrics
To validate our method experimentally, we used seven real-world datasets, all but one containing both prominent local and global structures (Table 1). We measured the embedding quality using two established metrics quantifying local and global structure preservation (Kobak and Berens, 2019):
  1. KNN
    The $k$-nearest neighbor recall (KNN) is the fraction of $k$-nearest neighbors in the high-dimensional data that are preserved as $k$-nearest neighbors in the low-dimensional embedding. We used $k=10$ throughout all experiments. The final metric is given as the average across all $n$ data points. KNN quantifies the preservation of local structure in the embedding.
  2. CPD
These two metrics rely solely on the high-dimensional data $X$ and its corresponding embedding $Y$ and do not make use of any metadata listed in Table 1.
This score ranges from 0 (embedding has the worst local and the worst global scores) to 1 (embedding has the best local and the best global scores), and allows direct comparison of embedding methods regarding their combined local and global structure preservation. Note that it only produces a relative score as it depends on the considered competitors.
@@ 5.2 Comparison methods
  • $t$-SNE (van der Maaten and Hinton, 2008) using the openTSNE implementation (Poličar et al., 2024) with the same initialization as in DREAMS (PCA normalized by the standard deviation of PC1), and UMAP (McInnes et al., 2018) as baselines for local structure preservation;
@@
  • TriMap (Amid and Warmuth, 2019), PacMAP (Wang et al., 2023), and PHATE (Moon et al., 2019) as a methods that strive to preserve both local and global structure;
@@
We always report means and uncertainties (often barely perceptible) across four random seeds. All experiments were conducted on a single Intel Xeon Gold 6226R CPU 2.90 GHz (16 cores, 32 threads), an NVIDIA RTX A6000 GPU (48 GB VRAM, CUDA 12.7), and 377 GB system RAM, running on a Linux environment. Only experiments including DREAMS-CNE or DREAMS-CNE-Decoder utilized the GPU. Without parallelization, the runtime of all experiments was about a week.
@@ 6.1 DREAMS successfully combines the strengths of $t$-SNE and PCA
Figure 2: Embeddings of the Tasic et al. dataset. a Spectrum of DREAMS embeddings for different values of $\lambda$. b Embeddings obtained by some of the competing methods. For all embeddings, see Figure S2.
@@
The quantitative metrics corroborated DREAMS ability to maintain both local and global structure (Figure 3). Across all seven datasets, $t$-SNE consistently achieved the highest KNN value, providing the best locally structured embedding. In contrast, PCA and MDS maintained global structure best, as reflected in their CPD values being the highest. For nearly all regularization parameters $\lambda\in[0,1]$ DREAMS yielded embeddings with better local or global structure preservation than its competitors. On the multi-scale datasets (all our datasets apart from MNIST), DREAMS with its default regularization strength ($\lambda=0.1$) simultaneously achieved KNN close to $t$-SNE’s and CPD close to PCA’s. Across all datasets, DREAMS preserved the local structure much better than StarMAP, which also combines neighbor embeddings with PCA. This is likely because DREAMS relies on $t$-SNE, which preserves local structure better than UMAP, that is used in StarMAP.
@@
UMAP generally performed poorly in our metrics, often preserving not only the local but also the global structure worse than $t$-SNE, as measured by the CPD metric. This is likely due to different default initializations of openTSNE and UMAP (PCA and Laplacian Eigenmaps, respectively). In contrast, TriMap often performed similarly well to StarMap, but still fell short compared to DREAMS in terms of local structure preservation. PHATE and PacMAP yielded worse CPD and worse KNN scores compared to DREAMS on all datasets.
@@
Only on the MNIST dataset DREAMS was not able to maintain high global and local structure at the same time, likely because this dataset does not have a prominent global structure in the first place (CPD values were comparatively low for all methods including PCA and MDS). Nevertheless, even here, DREAMS was on par or outperformed most other methods (Figure 3g).
@@
Using our aggregated local-global score, we could directly compare DREAMS to other methods along a single dimension (Table 2). The highest local-global score on every dataset was achieved either by DREAMS or by DREAMS-MDS, depending on whether PCA or MDS was capturing global structure more effectively in terms of the CPD metric. Following DREAMS, SQuadMDS-hybrid was consistently the next best method. It achieved scores comparable to DREAMS on the 1000 Genomes and MNIST datasets, but on all remaining datasets, its performance measured by the local-global score was at least $13\%$ lower than the best performing DREAMS variant. On the multi-scale datasets (all datasets apart from MNIST), all remaining other methods had scores at least $19\%$ lower than the best performing DREAMS variant.
@@
Table 2: Aggregated local-global score. For each dataset, the methods within 0.05 of the highest score are highlighted in bold.
Tasic
et al.
Macosko
et al.
Kanton
et al.
Wagner
et al.
Packer
et al.
1000
Genomes
MNIST
DREAMS 0.91 0.87 0.88 0.86 0.62 0.89 0.62
DREAMS-MDS 0.87 0.80 0.90 0.89 0.81 0.92 0.67
SQuadMDS-hybrid 0.71 0.61 0.77 0.76 0.63 0.88 0.61
StarMAP 0.68 0.56 0.66 0.59 0.38 0.66 0.26
PHATE 0.25 0.42 0.40 0.39 0.28 0.35 0.19
PaCMAP 0.45 0.35 0.42 0.39 0.35 0.51 0.20
TriMap 0.66 0.55 0.56 0.58 0.51 0.55 0.10
UMAP 0.26 0.46 0.26 0.19 0.39 0.31 0.19
$t$-SNE 0.60 0.50 0.60 0.63 0.50 0.73 0.61
PCA 0.50 0.50 0.48 0.46 0.24 0.43 0.35
MDS 0.48 0.47 0.52 0.51 0.51 0.55 0.50
@@ 6.2 DREAMS provides the best local-global spectrum
Next we compared the local-global embedding spectrum of DREAMS to the local-global spectra of existing methods and variants of DREAMS.
@@
Moreover, while DREAMS includes standard $t$-SNE and PCA/MDS as its corner cases for $\lambda=0$ and $\lambda=1$ (marked with stars in Figure 4), neither SQuadMDS-hybrid nor StarMAP could reach MDS and PCA, respectively, in their most global setting. Additionally, SQuadMDS-hybrid underperformed compared to $t$-SNE in its most local configuration (Figure 4a,b). This outcome is expected for StarMAP, as its objective continues to blend aspects of UMAP and PCA, even at its highest regularization strength. In the case of SQuadMDS-hybrid, the reason is likely the normalization of the $t$-SNE and MDS gradients before combining them.
Switching DREAMS’ neighbor-embedding backend from openTSNE to contrastive neighbor embeddings (CNE, Damrich et al. 2023) with the InfoNCE loss decreased the local structure preservation, directly proportionally to the difference in local structure preservation between $t$-SNE and its InfoNCE version InfoNC-$t$-SNE (Figure 4c). Using the CNE backend, we compared regularization using a precomputed PCA embedding (DREAMS-CNE) with linear decoding regularization (DREAMS-CNE-Decoder), and observed only marginal improvements with the decoder approach (Figure S1). The MNIST dataset was the only one where the decoder approach performed much better in terms of our metrics, but this did not translate into visible improvements of the embedding structure. Although the decoder could reconstruct the PCA embedding and enhance global structure, it introduces additional randomness and computational complexity. Moreover, with a precomputed global embedding, we are more flexible to adjust the global layout e.g. by switching from PCA to MDS. For these reasons and especially due to the better local structure of openTSNE than InfoNC-$t$-SNE, we prefer the simple regularization using a precomputed embedding as in the openTSNE-based DREAMS variants.
Figure 4: Trade-offs between local and global structure preservation in different methods. Spearman correlation of pairwise distances is plotted against kkNN recall. Panels a and c show results on the Tasic et al. dataset, while panel b is based on the Packer et al. dataset. a: Performance of DREAMS compared with other local-global spectra. b: Comparison of DREAMS-MDS and SQuadMDS-hybrid. c: Performance of different DREAMS variants.
@@ 7 Discussion
In this work, we introduced DREAMS, a dimensionality reduction method that adds a regularization term to the $t$-SNE objective. This regularization term penalizes embedding points that deviate from their corresponding PCA positions and thereby encourages global structure preservation throughout the optimization process. Our approach addresses a critical shortcoming in conventional dimensionality reduction techniques, which often prioritize one structural scale over the other. Through this simple yet effective regularization term and a tunable hyperparameter $\lambda$, DREAMS allows for a continuous spectrum of embeddings that transition smoothly from the locally faithful structure of $t$-SNE to the globally coherent and interpretable structure of PCA. Furthermore, with its default regularization strength ($\lambda=0.1$), DREAMS provides an embedding that successfully balances local and global structure preservation, outperforming competing methods.
@@
Across seven real-world datasets, DREAMS or DREAMS-MDS with default $\lambda=0.1$ consistently achieved the best performance in terms of the local-global score. This highlights its robustness and wide applicability, especially to hierarchical datasets with multiple inherent structure scales. While multiscale structure is common in real-world datasets, DREAMS is less useful in datasets without it, such as the MNIST dataset, as there is less global structure to preserve. Moreover, DREAMS introduces an additional hyperparameter, the regularization strength $\lambda$, whose optimal value can depend on the specific dataset and the intended use. Consequently, different datasets may require tuning $\lambda$ to achieve the optimal balance between local and global structure preservation. That said, we found that $\lambda=0.1$ worked well across the datasets.
@@
Because the regularization enforces alignment with a predefined structure, any biases or limitations in the PCA embedding would propagate into the final embedding. As demonstrated in cases where the MDS embedding captures global structure more effectively than PCA, sometimes it can be beneficial to regularize towards the MDS embedding, as in DREAMS-MDS. Thanks to its flexible and simple design, DREAMS accommodates such substitutions, making it adaptable to the specific characteristics of different datasets.
@@
Our alternative implementation DREAMS-CNE-Decoder performed worse than the default DREAMS due to the lower local quality of sampling-based InfoNC-$t$-SNE. Therefore we prefer our openTSNE-based implementation, even though the PyTorch implementation may be preferable in some use cases.
@@
In conclusion, DREAMS successfully combines the local structure preservation strength of $t$-SNE with the global structure of PCA. This makes DREAMS’ visualizations of high-dimensional data more faithful and interpretable, particularly for datasets with hierarchical structure.

### v2

@@ 4.2 Linear decoding regularization
$\widetilde{W}=\underset{W\in\mathbb{R}^{m\times d}}{\arg\min}\|X-XWW^{\top}\|_{F}^{2}\text{ subject to }W^{\top}W=I_{d}.$ (4)
Since the PCA embedding is given by $\widetilde{Y}=X\widetilde{W}$, the DREAMS loss with linear decoding regularizer becomes
$\mathcal{L}(Y,D)=(1-\lambda)\mathcal{L}_{t\text{-SNE}}+\lambda\left\|X-(YD^{\top}+b)\right\|_{F}^{2},$
with $D\in\mathbb{R}^{m\times d}$ being a trainable linear decoder and $b\in\mathbb{R}^{m}$ a trainable bias term that is added row-wise and allows to handle uncentered embeddings. In this setup, by minimizing the reconstruction error, the decoder is responsible for the global structure preservation by pushing the embedding towards the PCA structure that has the minimal reconstruction error. If $D$ were constrained to have orthogonal columns ($D^{\top}D=I_{d}$), the optimum of the regularization term would be $D=\widetilde{W}$ (Plaut, 2018; Nazari et al., 2023). Although we do not explicitly enforce orthogonality, we observed that the learned linear decoder $D$ naturally tends to have approximately orthogonal columns.
@@
For comparison, we also implemented a version of DREAMS using the CNE backend with precomputed PCA regularization as in Section 4.1. We will refer to this version as DREAMS-CNE. Here the gradient with respect to $Y$ was computed using autodifferentiation, including the $\|Y\|_{F}$ contribution to $\alpha$.
@@ 5 Experimental setup
5.1 Datasets and performance metrics
To validate our method experimentally, we used eleven real-world datasets, all but one containing both prominent local and global structures (Table 1). We measured the embedding quality using two established metrics quantifying local and global structure preservation (Kobak and Berens, 2019):
  1. KNN
    The $k$-nearest neighbor recall (KNN) is the fraction of $k$-nearest neighbors in the high-dimensional data that are preserved as $k$-nearest neighbors in the low-dimensional embedding. We used $k=10$ throughout all experiments (for results with different choices of $k$, see Figure S5). The final metric is given as the average across all $n$ data points. KNN quantifies the preservation of local structure in the embedding.
  2. CPD
This score ranges from 0 (embedding has the worst local and the worst global scores) to 1 (embedding has the best local and the best global scores) and allows direct comparison of embedding methods regarding their combined local and global structure preservation. Note that it only produces a relative score as it depends on the considered competitors. For an intuitive understanding of the local-global score, see Figure S4.
These metrics rely solely on the high-dimensional data $X$ and its corresponding embedding $Y$ and do not make use of any metadata listed in Table 1. In contrast, metrics like classification accuracy, or class separation measured via the Silhouette score, or agreement between clusters and classes measured via the adjusted Rand index, all require class labels.
Table 1: Datasets used in our experiments with a description of their local and global structure. For MNIST and Fashion MNIST, we only used the training set.
@@ 2. CPD
Mammoth 3D Mammoth skeleton overall body shape bone structure 50 000
Satellite satellite image crops soil color 6 soil types 6 435
FMNIST images of fashion items shoes/bags/garments 10 classes 60 000
@@
CIFAR10 vehicle and animal images vehicles vs animals 10 classes 60 000
@@ 5.2 Comparison methods
  • $t$-SNE (van der Maaten and Hinton, 2008) using the openTSNE implementation (Poličar et al., 2024) and UMAP (McInnes et al., 2018) as baselines for local structure preservation;
@@
  • TriMap (Amid and Warmuth, 2019), PacMAP (Wang et al., 2023), and PHATE (Moon et al., 2019) as methods that strive to preserve both local and global structure;
@@
We always report means and uncertainties (often barely perceptible) across four random seeds. All experiments were conducted on a single Intel Xeon Gold 6226R CPU 2.90 GHz (16 cores, 32 threads), an NVIDIA RTX A6000 GPU (48 GB VRAM, CUDA 12.7), and 377 GB system RAM, running on a Linux environment. Only experiments including DREAMS-CNE or DREAMS-CNE-Decoder utilized the GPU. Without parallelization, the runtime of all experiments was about a week. For a runtime analysis of DREAMS and its competitors, see Figure S2.
@@ 6.1 DREAMS successfully combines the strengths of $t$-SNE and PCA
Figure 2: Embeddings of the Tasic et al. dataset. a: Spectrum of DREAMS embeddings for different values of regularization strength $\lambda$. b–e: Embeddings obtained by some of the competing methods. For all embeddings, see Figure S6.
@@
The quantitative metrics corroborated the ability of DREAMS to maintain both local and global structure (Figure 3). On all datasets but Mammoth, $t$-SNE achieved the highest KNN value, providing the best locally structured embedding. In contrast, PCA and MDS maintained global structure best, as reflected in their CPD values being the highest. For nearly all regularization parameters $\lambda\in[0,1]$ DREAMS yielded embeddings with better local or global structure preservation than its competitors. On most datasets, DREAMS with its default regularization strength ($\lambda=0.15$) simultaneously achieved KNN close to $t$-SNE’s and CPD close to PCA’s. Across all datasets, DREAMS preserved the local structure much better than StarMAP, which also combines neighbor embeddings with PCA. This is likely because DREAMS relies on $t$-SNE, which preserves local structure better than UMAP, which is used in StarMAP.
@@
UMAP generally performed poorly in our metrics, often preserving not only the local but also the global structure worse than $t$-SNE, as measured by the CPD metric. This may be due to different default initializations of openTSNE and UMAP (PCA and Laplacian Eigenmaps, respectively). That said, note that UMAP is known to perform well in clustering-based metrics (Espadoto et al., 2019; Xiang et al., 2021; Lause et al., 2024), which we do not use for evaluation in this work.
TriMap performed similarly well to StarMap (note that both use PCA as initialization), but still fell short compared to DREAMS in terms of local structure preservation. PHATE and PacMAP yielded worse CPD and worse KNN scores compared to DREAMS on all datasets.
@@
On the MNIST dataset DREAMS was not able to maintain high global and local structure at the same time, likely because this dataset does not have a prominent global structure in the first place (CPD values were comparatively low for all methods including PCA and MDS). Similarly, on the pixel-valued CIFAR10 dataset, DREAMS incurred a clear trade-off between local and global structure. On this challenging dataset all methods severely struggled (Figure S16). Nevertheless, even in these two cases, DREAMS was on par or outperformed most other methods (Figure 3j,k).
@@
Using our aggregated local-global score, we could directly compare DREAMS to other methods along a single dimension (Table 2). The highest local-global score on every dataset was achieved by either DREAMS or DREAMS-MDS, depending on whether PCA or MDS was capturing global structure more effectively in terms of the CPD metric. Following DREAMS, SQuadMDS-hybrid was consistently the next best method. It achieved scores comparable to DREAMS on three datasets, but on all remaining datasets, its score was lower by at least 0.06 than the best performing DREAMS variant. On the datasets where DREAMS achieved a good balance between local and global performance (all but MNIST and CIFAR10), all other methods had scores lower by at least 0.09 than the best performing DREAMS variant.
@@
Table 2: Aggregated local-global score. For each dataset, the methods within 0.05 of the highest score are highlighted in bold. The order of the datasets is as in Table 1.
Tas Mac Kan Wag Pac 1kG Mam Sat FMN MNI C10
DREAMS 0.88 0.85 0.88 0.84 0.62 0.89 0.95 0.95 0.83 0.61 0.66
DREAMS-MDS 0.87 0.78 0.90 0.90 0.83 0.93 0.95 0.94 0.87 0.64 0.62
SQuadMDS-hybrid 0.71 0.61 0.77 0.76 0.64 0.88 0.94 0.84 0.81 0.61 0.58
StarMAP 0.68 0.57 0.66 0.59 0.39 0.66 0.68 0.70 0.48 0.26 0.16
PHATE 0.25 0.42 0.40 0.39 0.29 0.35 0.01 0.47 0.14 0.19 0.03
PaCMAP 0.45 0.34 0.42 0.39 0.36 0.51 0.70 0.35 0.12 0.20 0.15
TriMap 0.66 0.55 0.56 0.58 0.51 0.55 0.68 0.58 0.24 0.10 0.25
UMAP 0.26 0.46 0.26 0.19 0.40 0.31 0.68 0.39 0.18 0.19 0.16
$t$-SNE 0.57 0.50 0.60 0.61 0.50 0.64 0.86 0.50 0.55 0.61 0.56
PCA 0.50 0.50 0.48 0.46 0.25 0.43 0.60 0.48 0.44 0.35 0.32
MDS 0.48 0.47 0.52 0.51 0.51 0.55 0.50 0.54 0.51 0.50 0.50
@@ 6.2 DREAMS provides the best local-global spectrum
Next, we compared the local-global embedding spectrum of DREAMS to the local-global spectra of existing methods and variants of DREAMS.
@@
Moreover, while DREAMS includes standard $t$-SNE and PCA/MDS as its corner cases for $\lambda=0$ and $\lambda=1$ (marked with stars in Figure 4), neither SQuadMDS-hybrid nor StarMAP could reach MDS and PCA, respectively, in their most global setting. Additionally, SQuadMDS-hybrid underperformed compared to $t$-SNE in its most local configuration (Figure 4a,b). This outcome is expected for StarMAP, as its objective continues to blend aspects of UMAP and PCA even at its highest regularization strength. For SQuadMDS-hybrid, the reason is likely the normalization of the $t$-SNE and MDS gradients before combining them.
Switching the neighbor-embedding backend of DREAMS from openTSNE to contrastive neighbor embeddings (CNE; Damrich et al., 2023) with the InfoNCE loss decreased the KNN score, proportionally to the difference in KNN score between $t$-SNE and its InfoNCE version InfoNC-$t$-SNE (Figure 4c). Using the CNE backend, we compared regularization using a precomputed PCA embedding (DREAMS-CNE) with linear decoding regularization (DREAMS-CNE-Decoder) and observed only marginal improvements with the decoder approach (Figure S3). The MNIST dataset was the only one where the decoder approach performed much better in terms of our metrics, but this did not translate into visible improvements of the embedding structure. Although the decoder could reconstruct the PCA embedding and enhance global structure, it introduces additional randomness and computational complexity. Moreover, with a precomputed global embedding, we are more flexible to adjust the global layout, e.g., by switching from PCA to MDS. For these reasons and especially due to the better local structure of openTSNE than InfoNC-$t$-SNE, we prefer the simple regularization using a precomputed embedding as in the openTSNE-based DREAMS variants.
Figure 4: Trade-offs between local and global structure preservation in different methods. Spearman correlation of pairwise distances is plotted against kkNN recall. a: Performance of DREAMS compared with other local-global spectra. b: Comparison of DREAMS-MDS and SQuadMDS-hybrid. c: Performance of different DREAMS variants. Panels a and c show results on the Tasic et al. dataset, while panel b is based on the Packer et al. dataset.
@@ 7 Discussion
In this work, we introduced DREAMS, a dimensionality reduction method that adds a regularization term to the $t$-SNE objective. This regularization term penalizes embedding points that deviate from their PCA positions and thereby encourages global structure preservation throughout the optimization process. Our approach addresses a critical shortcoming in conventional dimensionality reduction techniques, which often prioritize one structural scale over the other. Through this simple yet effective regularization term and a tunable hyperparameter $\lambda$, DREAMS allows for a continuous spectrum of embeddings that transition smoothly from the locally faithful structure of $t$-SNE to the globally coherent and interpretable structure of PCA. Furthermore, with its default regularization strength ($\lambda=0.15$), DREAMS provides an embedding that successfully balances local and global structure preservation, outperforming competing methods.
@@
Across eleven real-world datasets, DREAMS or DREAMS-MDS with default $\lambda=0.15$ consistently achieved the best performance in terms of the local-global score. This highlights its robustness and wide applicability, especially to hierarchical datasets with multiple inherent structure scales. While multiscale structure is common in real-world datasets, DREAMS is less useful in datasets without it, such as the MNIST dataset, as there is less global structure to preserve. In particular, we do not have theoretical performance guarantees ensuring that DREAMS will balance local and global structure well. The additional time to compute the regularizer and its gradient led to DREAMS having slower runtime than most competitors (Figure S2). Moreover, DREAMS introduces an additional hyperparameter, the regularization strength $\lambda$, whose optimal value can depend on the specific dataset and the intended use. Consequently, different datasets may require tuning $\lambda$ to achieve the optimal balance between local and global structure preservation. That said, we found that $\lambda=0.15$ worked well across all datasets.
@@
Because the regularization enforces alignment with a predefined structure, any biases or limitations in the PCA embedding would propagate into the final embedding. As demonstrated in cases where the MDS embedding captures global structure more effectively than PCA, it can sometimes be beneficial to regularize towards the MDS embedding, as in DREAMS-MDS. Thanks to its flexible and simple design, DREAMS accommodates such substitutions, making it adaptable to the specific characteristics of different datasets.
@@
Our alternative implementation DREAMS-CNE-Decoder performed worse than the default DREAMS due to the lower local quality of sampling-based InfoNC-$t$-SNE. Therefore, we prefer our openTSNE-based implementation, even though the PyTorch implementation may be preferable in some use cases.
@@
In conclusion, DREAMS successfully combines the local structure preservation strength of $t$-SNE with the global structure of PCA. This makes visualizations of high-dimensional data more faithful and interpretable, particularly for datasets with hierarchical structure.